# Average Winding Number for Determinantal Curves associated to 2-Matrix Models in the Class AIII

### v1

@@ 1 Introduction
The discovery of the quantum Hall effect in 1980 [1] revealed that, as the strength of an external magnetic field increases, the Hall conductance of a two-dimensional electron gas forms quantized plateaux. These plateaux persist even in the presence of impurities or lattice imperfections, indicating a remarkable robustness of the quantization to disorder.
This phenomenon initiated the modern study of topological phases of matter, where insulating bulk behaviour coexists with conducting edge states. The systematic classification of such phases was achieved in the seminal work [2] and relies on the Altland-Zirnbauer classification [3], sometimes referred to as the Tenfold Way [4], which organises spectrally gapped Hamiltonians according to the Cartan classification of Riemannian symmetric spaces.
More recently, non-Hermitian Hamiltonians have attracted interest as effective models for open quantum systems where dissipation may occur. In that context the interplay between topology and spectral structure plays a central role, see [5] for a recent review.
Through the bulk-boundary correspondence (see for example [6, 7]), edge phenomena observed in systems with open boundary conditions may be understood by studying the same system with closed boundary conditions often referred to as the bulk. The corresponding bulk Hamiltonian has the aforementioned discrete translational lattice invariance. Its Fourier transform may be parametrized by $p\in\mathrm{S}^{1}$ on the unit circle when considering the one-dimensional setting. Furthermore its eigenvectors form a complex vector bundle over $\mathrm{S}^{1}$. Non-triviality for the latter corresponds to the existence of protected edge states in the original system.
@@
Among the ten Altland-Zirnbauer symmetry classes, the chiral class AIII exhibits a particularly rich structure in odd dimensions due to the presence of an energy gap at the Fermi level. It corresponds to quantum systems that are neither time-reversal invariant nor particle-hole invariant yet remain symmetric under the composition of time-reversal and particle-hole conjugation, see [8, Section E]. In such models, the determinant of the off-diagonal block of the Hamiltonian defines a map from the circle $\mathrm{S}^{1}$ to the punctured complex plane $\mathbb{C}^{*}$ whose winding number around the origin serves as a characteristic integer invariant.
As noted in [9, 10], this invariant given by the determinantal curve remains meaningful under random perturbations even though averaging may smear out the spectral gap. Indeed, each realization of the random Hamiltonian retains a finite gap whose typical length scales with the local mean level spacing.
@@
Our analysis extends the computation of the partition function associated with the random winding number, previously derived for the Ginibre Unitary Ensemble (GinUE) additive two-matrix model [9, 11, 10], to a broader class of non-Hermitian additive two-matrix models for which exact formulas can still be obtained, allowing us to extract large-dimension asymptotics for the average winding number.
Section 2 introduces the model and the associated constructions including the winding number, the partition function and Pólya ensembles of multiplicative type [12, 13, 14]. In this section, we also summarise our main results: Theorem 2.8 provides an exact expression valid for finite size $N$ and for all Pólya two-matrix models whereas Theorem 2.9 gives the asymptotic expansion of the mean winding number for the Muttalib-Borodin ensemble [15] of Laguerre type. As a comparison, 2.14 confirms that the result of [9, 11] for the GinUE additive 2-matrix model extends to more general matrix-valued Gaussian random fields. Proofs of the finite and large-$N$ results are presented in Sections 3 and 4, respectively, and the auxiliary integral identities used therein are collected in the appendices. We conclude in Section 5.
@@ 2.1 Winding number of a determinantal curve
where $(\cdot)^{\dagger}$ denotes the Hermitian conjugate and we recall that $p$ denotes the quasi-momentum varying over the Brillouin zone, which is isomorphic to the unit circle. The matrix-valued random field $K$ is assumed to be analytic entry-wise over $\rm{S}^{1}$.
@@
In a naive approach, the determinantal curve $p\mapsto\det\left(H\left(p\right)\right)$ would not yield any significant topological information since $H$ is Hermitian and its determinant takes values in the homotopically trivial space $\mathbb{R}_{+}$. Following [16], one may instead consider the determinantal curve over $\mathrm{S}^{1}$ associated with the sub-block $K$. We note that the eigenvalues of $K$ are not necessarily periodic and some might be permuted, see Fig 1.
Figure 1: Eigenvalue flow of $K(p)=(\frac{1}{2i}(p-p^{-1})+p)K_{1}+(3-\frac{1}{10}(p+p^{-1})+p^{-2})K_{2}$ where the two source random matrices $K_{1}$ and $K_{2}$ are independently drawn from the $\mathrm{GinUE}(5)$ and its associated determinantal curve. The black dots indicate the $5$ eigenvalues of $K(0)$ and $\det(K(0))$ while the red arrows mark the points where $\mathrm{Arg}(p)=\pi$. The winding number around the origin of the determinantal curve equals +1+1. The eigenvalue flow is not $2\pi$-periodic: as $p$ winds once around $\mathrm{S}^{1}$, one eigenvalue traces a closed orbit whereas the remaining four form a 44-cycle. As a result, the flow is $8\pi$-periodic.
@@
$\mathrm{Wind}_{N}\left(\mathscr{C},0\right)=\frac{1}{2\pi i}\oint_{\rm{S}^{1}}w\left(p\right)\mathrm{d}p,$ (2.3)
@@
This quantity will be our object of interest and we will study its $m$-point correlation function with $m\geq 1$ and $N\geq 1$. The arguments of $\mathrm{Wind}_{N}$ shall highlight the chosen curve $\mathscr{C}$ and the point it winds around. For short we denote the winding number by $\mathrm{Wind}_{N}=\mathrm{Wind}_{N}\left(\mathscr{C},0\right)$.
To investigate the distribution of that random winding number, a natural quantity to consider is the expectation of its $m$th moment as it opens the way for a possible Central Limit Theorem (CLT), i.e.,
@@
As shown above, the integration of $\mathrm{C}^{\left(N\right)}_{m}$ over $\mathbb{T}^{m}$ gives the $m$th moment of $\mathrm{Wind}_{N}$.
@@ 2.2 Additive 2-matrix models
The 2-matrix model with both random matrices are independently drawn from the GinUE, which is defined by the requirement that all entries are independent standard complex Gaussian random variables and was first introduced in [17], has been studied in detail recently in [11]. A recent survey may be found at [18].
@@
In the special case where the source matrices are drawn from the GinUE of size $N$ the result for the partition function is quite simple and was derived in [11, Sec. IV]:
@@
This result highlights an elegant and algebraically rich structure for the partition function associated to the random winding number.
We aim to expand this result to a larger class of random matrix ensembles and to include complex conjugates in the partition function (2.6) in order to allow for the possible capture of real and imaginary parts of the winding number density (2.4). This class of random matrices is called multiplicative Pólya ensembles on the complex general linear group $\mathrm{Gl}_{N}(\mathbb{C})$ and will be reviewed in the next subsection.
@@ 2.3 Pólya ensembles of multiplicative type on $\mathrm{Gl}_{N}(\mathbb{C})$
The cornerstone of the computation in [11] was to reduce the 2-matrix model to a single matrix model which consists of the product of the first matrix of the model multiplied by the inverse of the second one. This reduction is what ultimately makes the closed-form result (2.12) possible. To generalize this idea, we need a class of random matrices that is closed under product and inverse while retaining suitable integrable structure.
Such a class has been constructed [13, 14] in the past decade and is referred to as Pólya Ensembles on $\mathrm{Gl}_{N}\left(\mathbb{C}\right)$. It has many convenient analytical and algebraic properties, for instance both its collections of eigenvalues and of singular values form determinantal point processes. We will give a brief introduction, an in-depth treatment can be found in [12, 13, 14].
@@
A more complete treatment of the properties of the Mellin transform can be found in [13, Sec. 2].
@@
The second level $\mathrm{PF}_{2}$ also admits a complete characterization: a continuous and integrable function $f$ on $\mathbb{R}$ is in $\mathrm{PF}_{2}$ if and only if $f$ is positive and log-concave. A proof may be found in Ref. [19, Section 11]. We also signal that a complete, albeit more involved, characterization of $\mathrm{PF}_{3}$ is available in [20].
Sometimes coined Polynomial Ensembles of Derivative Type in the literature, Pólya ensembles on $\mathrm{Gl}_{N}\left(\mathbb{C}\right)$ are families of matrix-valued random variables parametrized by a Pólya frequency function $\omega$ which are included in polynomial ensembles firstly introduced in [21], themselves included in the larger family of bi-unitary invariant (sometimes called isotropic) random matrices [13, 14]. The bi-unitary invariance means that their probability measure on $\mathrm{Gl}_{N}\left(\mathbb{C}\right)$ is invariant under the adjoint action of the unitary group $\mathrm{U}(N)$ in the following sense
@@
This property implies that the associated joint probability density function (jpdf) of the complex eigenvalues denoted by $f_{\mathrm{EV}}$ is rotationally-invariant. It also implies that the singular values are independent from their singular vectors, each of the latter being Haar distributed on the unit sphere of $\mathbb{C}^{N}$. It happens that the distribution of singular values looks simpler when one squares them, the corresponding squared singular values jpdf will be denoted by $f_{\mathrm{SV}}$.
We say that a function $\omega\in\mathrm{L}^{1,N-1}_{\left[1,N\right]}\left(\mathbb{R}_{+}\right)$ gives rise to a Pólya Ensemble on ${\rm Gl}_{N}\left(\mathbb{C}\right)$ as in [14, Eq. (2.17)] (equivalently that $\omega$ is a Pólya weight), if and only if the logarithmic deformation
@@
The normalization constant corresponds to the first term of
@@
where $\psi$ is a continuous convex function on $\mathbb{R}$. Additionally, logarithmic deformations of Pólya weights in $\mathrm{PF}_{2}$ decay at least exponentially near infinity, see for example Lemma  A.1 in [22].
@@
$\check{\omega}(x)=\frac{1}{\mathcal{M}[\omega](N)x^{N+1}}\omega\left(\frac{1}{x}\right).$ (2.31)
Proposition 2.3 (see Eq. (3.10), Rem. 3.5 in [13])
@@
The previous proposition is crucial in ensuing analysis and allows us to explicitly reduce the 2-matrix model to a 1-matrix model. For a given Pólya weight $\omega$ we define the self-inverse-convoluted Pólya weight by
@@
A central non-Gaussian example is the Muttalib-Borodin ensemble of Laguerre type [15] closely connected to quantum transport models via the Dorokhov-Mello-Pereyra-Kumar (DMPK) equation [23, 24]. The corresponding choice of the Pólya weight, after re-labelling to avoid confusion with our own notation, is
@@
Special choices of parameters recover familiar ensembles: setting $\alpha=1$, $\delta=0$ and $\gamma=1$ yields the GinUE while $\alpha=1$, $\delta\neq 0$ and $\gamma=1$ corresponds to the induced GinUE (see for example  [25]). The Muttalib-Borodin ensemble of Laguerre type has been studied extensively in [15, Sec. 4] and the corresponding jpdf of its complex eigenvalues is
@@
$\int_{\mathbb{C}}|z|^{j}\omega(|z|^{2})\mathrm{d}^{2}z<\infty$ (2.43)
and it has support $\widetilde{\beta}_{1},\widetilde{\beta}_{2}]\subset[0,\infty[$. Let $m$ be a strictly positive integer, $(\alpha_{\ell})$ be $m$ complex numbers and $(\beta_{\ell})$ be $m$ pairwise distinct complex numbers which are not equal to $\widetilde{\beta}_{1}$ and $\widetilde{\beta}_{2}$. Then, the following holds
@@
On the complement $\Omega^{\complement}$ there exists at least one index pair $(j,\ell)$ such that $\left|z_{j}-\beta_{\ell}\right|<r$. The closed disc $\overline{\mathrm{D}}(\beta_{j},r)$ is compact. Hence, all terms except $1/(z_{j}-\beta_{\ell})$ remain bounded.
Since $\widetilde{\omega}$ is in $\mathrm{PF}_{2}$, it is continuous and locally bounded on the interior of its support which may be $\mathbb{R}$, a half-line or an interval. Because the points $\beta_{\ell}$ are fixed and distinct, we can choose $r>0$ so that $\left|z\right|^{2}$ stays in a compact region where $\omega(\left|z\right|^{2})$ is bounded. Consequently, all factors in the integrand remain bounded on the compact disc $\overline{\mathrm{D}}(\beta_{j},r)$ except for the term $1/\left|z_{j}-\beta_{\ell}\right|$.
$(1+\left|z_{j}\right|)^{2(N-1)}\omega(|z_{j}|^{2})|z_{j}-\alpha_{\ell}|\prod_{k\neq\ell}|z_{j}-\alpha_{k}|^{-1}$ (2.49)
in the integral over $\overline{\mathrm{D}}(\beta_{j},r)$ can be bounded by a constant. Finally, we recall that
$\int_{\overline{\mathrm{D}}(\beta_{j},r)}\frac{\mathrm{d}^{2}z}{\left|z-\beta_{j}\right|}=\int_{\overline{\mathrm{D}}(0,r)}\frac{\mathrm{d}^{2}z}{\left|z\right|}=\int_{0}^{2\pi}\mathrm{d}\theta\int_{0}^{r}\frac{1}{\rho}\rho\,\mathrm{d}\rho=2\pi r<+\infty.$ (2.50)
Therefore, the original integral is finite. ∎
@@
$\text{for }\,1\leq j<\ell\leq m_{1}:a(q_{j})b(q_{l})\neq a(q_{l})b(q_{j}),$
$\text{for }\,1\leq j<\ell\leq m_{2}:a(\widetilde{q}_{j})b(\widetilde{q}_{l})\neq a(\widetilde{q}_{l})b(\widetilde{q}_{j}),$ (2.52)
$\text{for }\,1\leq j\leq m_{1},\,1\leq\ell\leq m_{2}:a(q_{j})b(\widetilde{q}_{l})\neq a(\widetilde{q}_{l})b(q_{j}),$
@@
where $\kappa$ has been introduced at (2.10). By construction, the random matrix $K_{1}^{-1}K_{2}$ is also distributed from a Pólya ensemble, this time with a Pólya weight given by $\check{\omega}_{1}\circledast\omega_{2}$ so we may rewrite the previous generalized partition function as

### v2

# Average Winding Number for Determinantal Curves associated with 2-Matrix Models in the Class AIII

@@ 1 Introduction
The discovery of the quantum Hall effect in 1980 [klitzingNewMethodHighAccuracy1980] revealed that, as the strength of an external magnetic field increases, the Hall conductance of a two-dimensional electron gas forms quantized plateaux. These plateaux persist even in the presence of impurities or lattice imperfections, indicating a remarkable robustness of the quantization to disorder.
This phenomenon initiated the modern study of topological phases of matter, where insulating bulk behaviour coexists with conducting edge states. The systematic classification of such phases was achieved in the seminal work [SchnyderClassificationTopologicalInsulators2008] and relies on the Altland-Zirnbauer classification [AltlandNonstandardSymmetryClasses1997], sometimes referred to as the Tenfold Way [BaezTenfoldWay2020], which organises spectrally gapped Hamiltonians according to the Cartan classification of Riemannian symmetric spaces.
More recently, non-Hermitian Hamiltonians have attracted interest as effective models for open quantum systems where dissipation may occur. In that context, the interplay between topology and spectral structure plays a central role, see [AshidaNonHermitianPhysics2020] for a recent review.
Through the bulk-boundary correspondence (see for example [GrafBulkEdgeCorrespondenceDisordered2018, SawadaBulkBoundaryCorrespondenceErgodic2024]), edge phenomena observed in systems with open boundary conditions may be understood by studying the same system with closed boundary conditions often referred to as the bulk. The corresponding bulk Hamiltonian has the aforementioned discrete translational lattice invariance. Its Fourier transform may be parametrized by $p\in\mathrm{S}^{1}$ on the unit circle when considering the one-dimensional setting. Furthermore, its eigenvectors form a complex vector bundle over $\mathrm{S}^{1}$. Non-triviality of the latter corresponds to the existence of protected edge states in the original system.
@@
Among the ten Altland-Zirnbauer symmetry classes, the chiral class AIII exhibits a particularly rich structure in odd dimensions due to the presence of an energy gap at the Fermi level. It corresponds to quantum systems that are neither time-reversal invariant nor particle-hole invariant yet remain symmetric under the composition of time-reversal and particle-hole conjugation, see [ChiuClassificationTopologicalQuantum2016, Section E]. In such models, the determinant of the off-diagonal block of the Hamiltonian defines a map from the circle $\mathrm{S}^{1}$ to the punctured complex plane $\mathbb{C}^{*}$ whose winding number around the origin serves as a characteristic integer invariant.
As noted in [braunWindingNumberStatistics2022, hahnUniversalCorrelations2024], this invariant given by the determinantal curve remains meaningful under random perturbations even though averaging may smear out the spectral gap. Indeed, each realization of the random Hamiltonian retains a finite gap whose typical length scales with the local mean level spacing.
@@
Our analysis extends the computation of the partition function associated with the random winding number, previously derived for the Ginibre Unitary Ensemble (GinUE) additive $2$-matrix model [braunWindingNumberStatistics2022, hahnWindingNumberStatistics2023, hahnUniversalCorrelations2024], to a broader class of non-Hermitian additive $2$-matrix models for which exact formulas can still be obtained, allowing us to extract large-dimension asymptotics for the average winding number.
Section 2 introduces the model and the associated constructions including the winding number, the partition function and Pólya ensembles of multiplicative type [kieburgExactRelationSingular2016, kieburgProductsRandomMatrices2019, forsterPolynomialEnsemblesPolya2021]. In this section, we also summarise our main results: Theorem 2.8 provides an exact expression valid for finite size $N$ and for all Pólya $2$-matrix models whereas Theorem 2.9 gives the asymptotic expansion of the mean winding number for the Muttalib-Borodin ensemble [borodinBiorthogonalEnsembles1998] of Laguerre type. As a comparison, 2.14 confirms that the result of [braunWindingNumberStatistics2022, hahnWindingNumberStatistics2023] for the GinUE additive 2-matrix model extends to more general matrix-valued Gaussian random fields. Proofs of the finite and large-$N$ results are presented in Sections 3 and 4, respectively, and the auxiliary integral identities used therein are collected in the appendices. We conclude in Section 5.
@@ 2.1 Winding number of a determinantal curve
where $(\cdot)^{\dagger}$ denotes the Hermitian conjugate and we recall that $p$ denotes the quasi-momentum varying over the Brillouin zone, which is isomorphic to the unit circle. The matrix-valued random field $K$ is assumed to be analytic entrywise over $\mathrm{S}^{1}$.
@@
In a naive approach, the determinantal curve $p\mapsto\det\left(H\left(p\right)\right)$ would not yield any significant topological information since $H$ is Hermitian and its determinant takes values in the homotopically trivial space $\mathbb{R}_{+}$. Following [MaffeiTopologicalCharacterizationChiral2018], one may instead consider the determinantal curve over $\mathrm{S}^{1}$ associated with the sub-block $K$. We note that the eigenvalues of $K$ are not necessarily periodic and some might be permuted, see Figure 1.
Figure 1: On the left: eigenvalue flow of $K(p)=(\frac{1}{2i}(p-p^{-1})+p)K_{1}+(3-\frac{1}{10}(p+p^{-1})+p^{-2})K_{2}$ where the two source random matrices $K_{1}$ and $K_{2}$ are independently drawn from the $\mathrm{GinUE}(5)$. On the right: its associated determinantal curve. The black dots indicate the $5$ eigenvalues of $K(0)$ and $\det(K(0))$ while the red arrows mark the points where $\mathrm{Arg}(p)=\pi$. The winding number around the origin of the determinantal curve equals +1+1. The eigenvalue flow is not $2\pi$-periodic: as $p$ winds once around $\mathrm{S}^{1}$, one eigenvalue traces a closed orbit whereas the remaining four form a 44-cycle. As a result, the flow is $8\pi$-periodic. Plots generated with Mathematica
@@
$\mathrm{Wind}_{N}\left(\mathscr{C},0\right)=\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}w\left(p\right)\mathrm{d}p,$ (2.3)
@@
This quantity will be our object of interest and we will study its $m$-point correlation functions with $m\geq 1$ and $N\geq 1$. The arguments of $\mathrm{Wind}_{N}$ shall highlight the chosen curve $\mathscr{C}$ and the point it winds around. For short we denote the winding number by $\mathrm{Wind}_{N}=\mathrm{Wind}_{N}\left(\mathscr{C},0\right)$.
To investigate the distribution of that random winding number, a natural quantity to consider is the expectation of its $m$th moment as it opens the way for a possible central limit theorem, i.e.,
@@
As shown above, the integrating $\mathrm{C}^{\left(N\right)}_{m}$ over $\mathbb{T}^{m}$ gives the $m$th moment of $\mathrm{Wind}_{N}$.
@@ 2.2 Additive 2-matrix models
The 2-matrix model for which both random matrices are independently drawn from the GinUE, which is defined by the requirement that all entries are independent standard complex Gaussian random variables and was first introduced in [ginibreStatisticalEnsemblesComplex1965], has been studied in detail recently in [hahnWindingNumberStatistics2023]. A recent survey may be found in [byunProgressStudyGinibre2025].
@@
In the special case where the source matrices are drawn from the GinUE of size $N$ the result for the partition function is quite simple and was derived in [hahnWindingNumberStatistics2023, Sec. IV]:
@@
This result highlights an elegant and algebraically rich structure for the partition function associated with the random winding number.
We aim to expand this result to a larger class of random matrix ensembles and to include complex conjugates in the partition function (2.6) in order to capture the real and imaginary parts of the winding number density (2.4). This class of random matrices is called multiplicative Pólya ensembles on the complex general linear group $\mathrm{GL}_{N}(\mathbb{C})$ and will be reviewed in the next subsection.
@@ 2.3 Pólya ensembles of multiplicative type on $\mathrm{GL}_{N}(\mathbb{C})$
The cornerstone of the computation in [hahnWindingNumberStatistics2023] was to reduce the 2-matrix model to a single matrix model which consists of the product of the first matrix with the inverse of the second one. This reduction is what ultimately makes the closed-form result (2.12) possible. To generalize this idea, we need a class of random matrices that is closed under product and inverse while retaining suitable integrable structure.
Such a class has been constructed [kieburgProductsRandomMatrices2019, forsterPolynomialEnsemblesPolya2021] in the past decade and is referred to as Pólya ensembles on $\mathrm{GL}_{N}\left(\mathbb{C}\right)$. It has many convenient analytical and algebraic properties, for instance both its collections of eigenvalues and of singular values form determinantal point processes. We will give a brief introduction, an in-depth treatment can be found in [kieburgExactRelationSingular2016, kieburgProductsRandomMatrices2019, forsterPolynomialEnsemblesPolya2021].
@@
A more complete treatment of the properties of the Mellin transform can be found in [kieburgProductsRandomMatrices2019, Sec. 2].
@@
The second level $\mathrm{PF}_{2}$ also admits a complete characterization: a continuous and integrable function $f$ on $\mathbb{R}$ is in $\mathrm{PF}_{2}$ if and only if $f$ is positive and log-concave. A proof may be found in Ref. [saumardLogconcavityStrongLogconcavity2014, Section 11]. We also signal that a complete, albeit more involved, characterization of $\mathrm{PF}_{3}$ is available in [WeinbergerCharacterizationPolyaFrequency1983].
Sometimes coined Polynomial Ensembles of Derivative Type in the literature, Pólya ensembles on $\mathrm{GL}_{N}\left(\mathbb{C}\right)$ are families of matrix-valued random variables parametrized by a Pólya frequency function $\omega$ which are included in polynomial ensembles firstly introduced in [kuijlaarsTransformationsPolynomialEnsembles2016], themselves included in the larger family of bi-unitary invariant (sometimes called isotropic) random matrices [kieburgProductsRandomMatrices2019, forsterPolynomialEnsemblesPolya2021]. The bi-unitary invariance means that their probability measure on $\mathrm{GL}_{N}\left(\mathbb{C}\right)$ is invariant under the adjoint action of the unitary group $\mathrm{U}(N)$ in the following sense
@@
This property implies that the associated joint probability density function (jpdf) of the complex eigenvalues denoted by $f_{\mathrm{EV}}$ is rotationally-invariant. It also implies that the singular values are independent of their singular vectors, each of the latter being Haar distributed on the unit sphere of $\mathbb{C}^{N}$. It happens that the distribution of singular values looks simpler when one squares them, the corresponding squared singular values jpdf will be denoted by $f_{\mathrm{SV}}$.
We say that a function $\omega\in\mathrm{L}^{1,N-1}_{\left[1,N\right]}\left(\mathbb{R}_{+}\right)$ gives rise to a Pólya Ensemble on ${\rm Gl}_{N}\left(\mathbb{C}\right)$ as in [forsterPolynomialEnsemblesPolya2021, Eq. (2.17)] (equivalently that $\omega$ is a Pólya weight), if and only if the logarithmic deformation
@@
The normalisation constants corresponds to the first term of
@@
where $\psi$ is a continuous convex function on $\mathbb{R}$. Additionally, logarithmic deformations of Pólya weights in $\mathrm{PF}_{2}$ decay at least exponentially near infinity, see for example Lemma  A.1 in [DumbgenMaximumLikelihoodEstimation2009].
@@
$\check{\omega}_{1}(x)=\frac{1}{\mathcal{M}[\omega_{1}](N)x^{N+1}}\omega_{1}\left(\frac{1}{x}\right).$ (2.31)
Proposition 2.3 (see Eq. (3.10), Rem. 3.5 in [kieburgProductsRandomMatrices2019])
@@
The previous proposition is crucial in the ensuing analysis and allows us to explicitly reduce the 2-matrix model to a 1-matrix model. For a given Pólya weight $\omega$ we define the self-inverse-convoluted Pólya weight by
@@
A central non-Gaussian example is the Muttalib-Borodin ensemble of Laguerre type [borodinBiorthogonalEnsembles1998] closely connected to quantum transport models via the Dorokhov-Mello-Pereyra-Kumar (DMPK) equation [dorokhovTransmissionCoefficientLocalization1982, melloMacroscopicApproachMultichannel1988]. The corresponding choice of the Pólya weight, after re-labelling to avoid confusion with our own notation, is
@@
Special choices of parameters recover familiar ensembles: setting $\alpha=1$, $\delta=0$ and $\gamma=1$ yields the GinUE while $\alpha=1$, $\delta\neq 0$ and $\gamma=1$ corresponds to the induced GinUE (see for example  [fischmannInducedGinibreEnsemble2012]). The Muttalib-Borodin ensemble of Laguerre type has been studied extensively in [borodinBiorthogonalEnsembles1998, Sec. 4] and the corresponding jpdf of its complex eigenvalues is
@@
$\int_{\mathbb{C}}|z|^{j}\omega(|z|^{2})\mathrm{d}^{2}z<\infty.$ (2.43)
Let $m$ be a strictly positive integer, $(\alpha_{\ell})$ be $m$ complex numbers and $(\beta_{\ell})$ be $m$ pairwise distinct non-zero complex numbers. Then, the following holds
@@
On the complement $\Omega^{\complement}$, there exists an index pair $(j,\ell)$ such that $\left|z_{j}-\beta_{\ell}\right|<r$. We work inside the compact closed disc
$\overline{\mathrm{D}}(\beta_{\ell},r)=\left\{z\in\mathbb{C}:\,\left|z-\beta_{\ell}\right|\leq r\right\}.$ (2.49)
Because $\omega$ is a Pólya weight, it is positive, continuous and integrable on $\mathbb{R}_{+}$. In particular, any singularity must arise at a boundary point say $t_{0}\in\mathbb{R}_{+}$: there exists $\alpha\in[0,1[$ and $c_{1}>0$ such that, on a neighbourhood of $t_{0}$
$\omega(t)\leq c_{1}\left|t-t_{0}\right|^{-\alpha}$
If $\left|\beta_{\ell}\right|^{2}\neq t_{0}$, we simply have $\alpha=0$ in what follows. Assume there exist an index $\ell$ such that $\left|\beta_{\ell}\right|^{2}=t_{0}$. Because $\beta_{\ell}$ is non-zero and the function $z\mapsto\left|z\right|^{2}$ has a non-vanishing gradient at $z=\beta_{\ell}$, for $r>0$ sufficiently small there exist constants $c_{2},c_{3}>0$ such that
$c_{2}\left|z-\beta_{\ell}\right|\leq\left|\left|z\right|^{2}-\left|\beta_{\ell}\right|^{2}\right|\leq c_{3}\left|z-\beta_{\ell}\right|.$
Consequently, on $\overline{\mathrm{D}}(\beta_{\ell},r)$ there exists $c_{4}>0$ such that
$\omega(\left|z\right|^{2})\leq c_{4}\left|z-\beta_{\ell}\right|^{-\alpha}.$
All factors in the integrand that remain away from their own singularities are bounded on this disc. In particular,
  • all denominators $\left|z_{j}-\beta_{k}\right|$ with $k\neq\ell$ stay uniformly bounded away from zero since the discs are chosen disjoints,
  • the Vandermonde term is bounded by
    $\left|\Delta_{N}(\mathbf{z})\right|^{2}\leq\prod_{j=1}^{N}(1+\left|z_{j}\right|)^{2(N-1)}.$
It follows that, for $z_{j}\in\overline{\mathrm{D}}(\beta_{\ell},r)$, the factor
$(1+\left|z_{j}\right|)^{2(N-1)}\omega(\left|z_{j}\right|^{2})\left|z_{j}-\alpha_{\ell}\right|\prod\limits_{k\neq\ell}\left|\frac{z_{j}-\alpha_{k}}{z_{j}-\beta_{k}}\right|$
is bounded, up to a constant, by $\left|z_{j}-\beta_{\ell}\right|^{-\alpha}$. Hence, there exists $c_{5}>0$ such that the full integrand on $\overline{\mathrm{D}}(\beta_{\ell},r)$ is dominated by
$c_{5}\left|z_{j}-\beta_{\ell}\right|^{-(1+\alpha)}.$
In polar coordinates around $\beta_{\ell}$, that is $z=\beta_{\ell}+\rho e^{i\theta}$, on this closed disc we have
$\int_{\overline{\mathrm{D}}(\beta_{\ell},r)}\left|z_{j}-\beta_{\ell}\right|^{-(1+\alpha)}\mathrm{d}^{2}z=\int_{0}^{2\pi}\mathrm{d}\theta\int_{0}^{r}\rho^{-(1+\alpha)}\rho\mathrm{d}\rho=2\pi\int_{0}^{r}\rho^{-\alpha}\mathrm{d}\rho<\infty,$
since $\alpha<1$. Therefore, the integral over $\Omega^{\complement}$ is finite which allows us to conclude.
@@
$\text{for }\,1\leq j<\ell\leq m_{1}:a(q_{j})b(q_{\ell})\neq a(q_{\ell})b(q_{j}),$
$\text{for }\,1\leq j<\ell\leq m_{2}:a(\widetilde{q}_{j})b(\widetilde{q}_{\ell})\neq a(\widetilde{q}_{\ell})b(\widetilde{q}_{j}),$ (2.51)
$\text{for }\,1\leq j\leq m_{1},\,1\leq\ell\leq m_{2}:a(q_{j})b(\widetilde{q}_{\ell})\neq a(\widetilde{q}_{\ell})b(q_{j}),$
@@
where $\kappa$ has been introduced at (2.10). By construction, the distribution of the random matrix $K_{1}^{-1}K_{2}$ is that of a Pólya ensemble, this time with a Pólya weight given by $\check{\omega}_{1}\circledast\omega_{2}$ so we may rewrite the previous generalized partition function as